For a particular combination of capital and labor we know that the marginal product of capital...

  1. For a particular combination of capital and labor we know that the marginal product of capital is 6 units of output and that the marginal rate of technical substitution is 3 units of capital per unit of labor. What is the marginal product of labor? Explain.
Answer #1

MRTS=MPL/MPK

MPK=Marginal product of capital

MPL=Marginal product of labour

MPL=MRTS*MPK

MPL=3*6=18

MPL=18